OpenStudy (anonymous):

Erin has developed a new lotion that will help children with sensitive skin. She invests $15,225 in equipment to manufacture and package her lotion. Each bottle costs $3.25 to manufacture and sells for $12. How many bottles of lotion must she make and sell before her business breaks even? a. 998 bottles b. 1740 bottles c. 174 bottles d. 1269 bottles HELLLPPP

OpenStudy (anonymous):

@yrelhan4

OpenStudy (jdoe0001):

hmm, what's breaking even?

OpenStudy (jdoe0001):

sounds like cutting thing in even pieces, like bread

OpenStudy (anonymous):

breaking even means like making a profit. lol

OpenStudy (jdoe0001):

well, does xxkerstie.bbyxx know that? I mean how can you find what you don't even know what it's

OpenStudy (anonymous):

The answer is B. Just take the manufacture's price of $3.25 and subtract it from $12, her cost. You get $8.75. Then just divide $15,225 by $8.75 and you get 1740 bottles.

OpenStudy (ybarrap):

Initial Cost: 15,225 Cost to make per bottle: 3.25 Selling Price per bottle: 12 Let X represent the number of bottles needed to break even: Profit: 12*X Expense: 15,225 + 3.25X Break Even Point = Profit - Expense = 0 12X - (15,225 + 3.25X)=0 X(12 - 3.25)= 15,225 X=\( \dfrac{15,225}{12-3.25}\), bottles needed to break even.